# FAIRGAME: a Framework for AI Agents Bias Recognition using Game Theory

### v2

@@ 4.2 Battle of sexes
Similarly to Fig. 2, Fig. 4 presents bar plots (with $95\%$ CI) summarizing the experimental results, over repeated experiments, under two conditions: when agents operate without knowledge of their opponent’s personality and when such information is provided. The data are reported for all LLM and languages evaluated in this study, and for all combinations of agents’ personalities. In this case, Llama 3 and Mistral Large show the highest internal variability, while Claude Sonnet and GPT-4o are very precise in their final output. Overall, the agents tend to cooperate to maximize the payoffs; however, if the personality if ’selfish’, cooperation drops dramatically and low payoffs are achieved. ’French’ agents are more cooperative than others, for all LLMs. The observed cooperation aligns with the equilibrium in coordination games like the Battle of the Sexes, where agents prioritize coordination over individual preferences to maximize collective payoffs. However, the sharp drop in cooperation with selfish personalities reflects the inherent difficulty in achieving equilibrium when agents prioritize individual payoffs over mutual benefit.
Figure 5: Evolution of normalized penalties (averaged over repeated experiments) over repeated rounds, for the Battle of sexes.
In Fig. 5, the outcomes generated by the LLM at each round were normalized to a scale from -1 to 1, where -1 corresponds to the maximum payoff (10), and 1 corresponds to the minimum payoff (0). Thus, these normalized values indicate cooperative rather than adversarial behavior between agents. Over rounds, all LLMs tend to converge to high payoffs, except for Llama, which 405B remains consistent throughout successive rounds around 0, suggesting that the cooperation level between the two agents does not improve over time. The convergence to high payoffs reflects the equilibrium-seeking behavior in coordination games like the Battle of the Sexes, where agents prioritize mutual benefit over individual preferences. The lack of improvement in Llama 3 405B’s cooperation suggests a rigid strategy that fails to adapt to repeated interactions, potentially missing opportunities to achieve higher collective payoffs through coordination.
@@ 4.3 Scoring system
    Sensitivity to Payoff ($S_{P}$): the model’s responsiveness to changes in incentives. We compute this by measuring the difference in behavior between the harsh ($H$) and mild ($M$) variants: $S_{P}=\frac{1}{Z_{S}}[\text{Mean}_{d}(|y^{(H)}_{d}-y^{(M)}_{d}|)]$, where $y^{(\cdot)}_{d}$ are the results for each round $d$, averaged over all other features.
@@
Fig. 6a shows such radar plot for the Prisoner’s dilemma. The higher the metric, the worse the LLM in a certain dimension; the area under the polygon gives immediate information about the overall performance. Mistral Large exhibits the highest variability across evaluation rounds and internal variability, whereas GPT-4o displays the highest sensitivity to payoffs. In contrast, Llama3 405B emerges as the model with the most stable overall behavior across the evaluated dimensions. Fig. 6b provides a comparative analysis of the LLMs in the Battle of Sexes game. Given that the LLM were not tested against multiple versions of the game featuring different payoff matrices, the metric $S_{P}$ was excluded from this evaluation.
Comparing Figs. 6a and b reveals that Mistral Large exhibits the greatest inconsistency across different languages, coupled with substantial internal variability, comparable to GPT-4o. Claude 3.5 demonstrates the highest variability across rounds. Lastly, although Llama3 405B shows notable internal variability, its behavior remains consistent across different languages and rounds; notably, this lower variability for Llama models (coupled with sometimes inconsistent results compared to other LLMs and predictions) was observed in other tasks [10], and appears to be a typical trait of the LLM.
@@ 4.4 Interpreting the results
To test the hypothesis, we first ask LLMs about their knowledge level on classic game theory scenarios. Results indicated that LLMs possess substantial familiarity with these games, including the optimal strategies, underlying mechanics, and associated sociological implications contrasting selfish and cooperative human behaviors.
To discern whether observed cooperative behaviors are a result of intrinsic predispositions or pre-existing knowledge, we modify the template file using different narratives for the game. For instance, we reframe the Prisoner’s Dilemma into a plane crash scenario, with survivors deciding whether to cooperate in collective hunting tasks. Instead of numerical payoffs, consequences were communicated qualitatively (e.g., "failure to cooperate results in starvation"). Despite these changes, cooperation levels remained consistently high, in line with Fig. 2. When explicitly queried, LLMs identified these disguised scenarios as variants of the Prisoner’s Dilemma, thereby complicating efforts to definitively attribute their behavior to either inherent biases or recognition of known game structures. Finally, we introduce distinct identities and personalities to the agents involved in these interactions to assess the impact on behavior. Our findings revealed notable behavioral shifts aligned with the assigned identities. Specifically, pairing archetypal figures such as Adolf Hitler, representing aggressive selfishness, and Mahatma Gandhi, symbolizing peaceful cooperation, resulted in predictable outcomes where the aggressive figure consistently opted for betrayal and the cooperative figure consistently opted to cooperate.
@@ 5 Conclusions
Identifying biases and predicting emerging behaviors in multi-LLM-agent interactions, using game-theoretic approaches, provide better interpretability and predictability to ensure fairness, operational efficiency, compliance with legal standards, and enhanced trust [5, 29, 46, 13]. Through quantification of interaction outcomes and their alignment with desired objectives using game-theoretic principles, FAIRGAME serves as a powerful instrument for assessing chatbot behavior, AI decision-making processes, and agent strategies across various contexts. Overall, our method quantifies the outcomes and discrepancies across LLM agents, in different games with varying characteristics. Our results also suggest that LLMs draw from their inner knowledge about games and characters involved, on top of the agnostic information contained in the description and payoff matrices, in order to choose their strategies. The scores computed in Sec. 6 are likely correlated with the degree of influence and bias provided by the training data, as well as with the tendency of LLMs to reduce statistical fluctuations at the cost of not evolving over rounds (as for Llama in the Game of sexes). Pairing quantitative results and qualitative observations about the expected outcomes of known games thus enables us to interpret the hidden biases of LLMs and to assess their emerging behaviors.
@@
FAIRGAME paves the way for a novel integration of LLMs and game theory, establishing a bidirectional relationship between them. Game theory provides the mathematical foundation for understanding strategic decision-making, interpreting and explaining how AI agents reason and make decisions. While, LLMs, as experimental tools for data-driven modeling of human decision-making, offer opportunities for empirical validation and exploration of complex interaction scenarios.
Future works may leverage FAIRGAME’s versatility and capabilities in numerous directions. First, additional biases can be investigated: on top of agents’ languages and personality, other factors can be included, such as nationality, gender, race, and age, so as to explore more nuanced characteristics and their influence on game outcomes and agent behavior. Moreover, future works can include agents’ communication, which is known to yield modified payoffs and strategies [33] and that may change the responses of agents possessing different languages or personality traits. Extending all evaluations to more than two agents would also enable the modeling of more complex interactions and strategies, such as coalition and commitment formations, and group reciprocity [48, 54, 49]. This would shed light onto how different combinations of agent traits influence group dynamics, identifying configurations that foster harmony and lead to improved collective performance for multi-agent systems, team formation strategies, cooperative AI, and the design of more socially intelligent artificial agents [18, 37].
FAIRGAME can also be extended to additional game-theoretic scenarios, even beyond "classical" games that are known to LLMs via training: these additional scenarios include incomplete information games, where players have partial or uncertain knowledge about the game state or other players’ strategies, simultaneous games, where players make decisions at the same time without knowing the actions of others, or sequential games, where players act in turns and can base their strategies on prior actions. Moreover, realistic game scenarios, where agents interact through real-world processes and their payoff are not pre-defined but inferred from the dynamics and outcomes of the actual use case, will allow us to predict plausible strategic outcomes. These inclusions will extend the scope of FAIRGAME and make it the preferred choice for integrating game-theory and LLMs and address advance game theoretic questions such as how biases in language or personality traits influence the emergence of cooperation or defection in evolutionary game dynamics, or what mechanisms can promote stable cooperation in multi-agent systems with diverse traits and incomplete information.
Moreover, considering the emerging paradigm represented by Agentic AI, where autonomous AI systems are designed to achieve complex goals with minimal human intervention [3], FAIRGAME can be extended to model interactions between AI-driven processes within organizations. These processes, representing agentic AI systems, influence and impact one another. Achieving optimal and productive cooperation among these systems is an aspirational target that requires careful consideration of personalities, cognitive behavioral traits, and environmental factors. In such scenarios, FAIRGAME can explore how diverse personalities, languages, and biases affect group dynamics, enabling the identification of optimal configurations that prioritize collective outcomes over individual interests. This extension is crucial for fostering effective coordination and cooperation among AI systems, enhancing adaptability, advanced decision-making capabilities, and self-sufficiency in dynamic environments. In sectors such as healthcare, finance, manufacturing, autonomous vehicles, cybersecurity, and smart cities, where AI systems must collaborate autonomously to achieve complex goals, understanding their interactions is essential for improving coordination, adaptability, and overall system performance.
@@
D.P. is supported by the European Union through the ERC INSPIRE grant (project number 101076926). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. T.A.H. is supported by EPSRC (grant EP/Y00857X/1).
FAIRGAME is open-source and accessible on Github [1], with examples of configuration files and templates employed in this work.

### v3

@@ 4.2 Battle of sexes
Similarly to Fig. 2, Fig. 4 presents bar plots (with $95\%$ CI) summarizing the experimental results, over repeated experiments, under two conditions: when agents operate without knowledge of their opponent’s personality and when such information is provided. The data is reported for all LLM and languages evaluated in this study, and for all combinations of agents’ personalities. In this case, Llama 3.1 405B and Mistral Large show the highest internal variability, while Claude Sonnet and GPT-4o are very precise in their final output. Overall, the agents tend to cooperate to maximize the payoffs; however, if the personality if ’selfish’, cooperation drops dramatically and low payoffs are achieved. ’French’ agents are more cooperative than others, for all LLMs. The observed cooperation aligns with the equilibrium in coordination games like the Battle of the Sexes, where agents prioritize coordination over individual preferences to maximize collective payoffs. However, the sharp drop in cooperation with selfish personalities reflects the inherent difficulty in achieving equilibrium when agents prioritize individual payoffs over mutual benefit.
Figure 5: Average evolution of coordination in strategy choices across repeated rounds for all experiments conducted in the Battle of the Sexes, shown separately for each LLM. As this is a coordination game, the plot examines whether the two LLMs select the same option in each round. A value of 1 indicates a mismatch in strategies (one selects Option A, the other Option B), reflecting coordination failure or defective behavior, while -1 indicates alignment in choices, reflecting successful coordination or cooperative behavior.
Figure 5 shows the evolution of strategy coordination across rounds in the Battle of Sexes game. All LLMs display a gradual improvement in coordination over time. Notably, in the first round, Mistral Large, Llama 3.1 405B Instruct, and GPT-4o tend to select opposing strategies, indicating a lack of coordination, whereas Claude 3.5 Sonnet demonstrates a higher level of alignment from the outset. By the end of the repeated interactions, Llama 3.1 405B Instruct remains largely uncoordinated on average, Mistral Large shows moderate misalignment, GPT-4o achieves moderate coordination, and Claude 3.5 approaches near-perfect coordination.
@@ 4.3 Scoring system
[…]
@@
Fig. 6a shows such a radar plot for the Prisoner’s dilemma. The higher the metric, the worse the LLM in a certain dimension; the area under the polygon gives immediate information about the overall performance. Mistral Large exhibits the highest variability across evaluation rounds and internal variability, whereas GPT-4o displays the highest sensitivity to payoffs. In contrast, Llama 3.1 405B emerges as the model with the most stable overall behavior across the evaluated dimensions. Fig. 6b provides a comparative analysis of the LLMs in the Battle of Sexes game. Given that the LLM were not tested against multiple versions of the game featuring different payoff matrices, the metric $S_{P}$ was excluded from this evaluation.
The scores computed in Sec. 6 are likely correlated with the degree of influence and bias provided by the training data, as well as with the tendency of LLMs to reduce statistical fluctuations at the cost of not evolving over rounds. Comparing Figs. 6a and b reveals that Mistral Large exhibits the greatest inconsistency across different languages, coupled with substantial internal variability, comparable to GPT-4o. Claude 3.5 demonstrates the highest variability across rounds. Lastly, although Llama 3.1 405B shows notable internal variability, its behavior remains consistent across different languages and rounds; notably, this lower variability for Llama models (coupled with sometimes inconsistent results compared to other LLMs and predictions) was observed in other tasks [11], and appears to be a typical trait of the LLM.
@@ 4.4 Interpreting the results
To test the hypothesis, we first asked LLMs about their knowledge level on classic game theory scenarios. Results indicated that LLMs possess substantial familiarity with these games, including the optimal strategies, underlying mechanics, and associated sociological implications contrasting selfish and cooperative human behaviors.
To discern whether observed cooperative behaviors are a result of intrinsic predispositions or pre-existing knowledge, we modified the template file using different narratives for the game. For instance, we reframed the Prisoner’s Dilemma into a plane crash scenario, with survivors deciding whether to cooperate in collective hunting tasks in an unhabited island. Instead of numerical payoffs, consequences were communicated qualitatively (e.g., "failure to cooperate results in starvation"). Despite these changes, cooperation levels remained consistently high, in line with Fig. 2. When explicitly queried, LLMs precisely identified these disguised scenarios as variants of the Prisoner’s Dilemma, thereby complicating efforts to definitively attribute their behavior to either inherent biases or recognition of known game structures. Finally, we run a second set of experiments introducing distinct identities and personalities to the agents involved in these interactions, so as to assess the impact on behavior. Our findings revealed notable behavioral shifts aligned with the assigned identities. Specifically, pairing archetypal figures such as Adolf Hitler, representing aggressive selfishness, and Mahatma Gandhi, symbolizing peaceful cooperation, resulted in predictable outcomes where the aggressive figure consistently opted for betrayal and the cooperative figure consistently opted to cooperate; this hints to LLMs using prior knowledge on top of the information encoded in the payoff matrices.
@@ 5 Conclusions
FAIRGAME provides a novel integration of LLMs and game theory, establishing a bidirectional relationship between them. Game theory provides the mathematical foundation for understanding strategic decision-making, interpreting and explaining how AI agents reason and make decisions; LLMs, as experimental tools for data-driven modeling of human decision-making, offer opportunities for empirical validation and exploration of complex interaction scenarios.
Applying game-theoretic approaches to multi-LLM interactions uncovers biases and emergent behaviors, improving interpretability, fairness, efficiency, legal compliance and trust [6, 45, 13]. The framework quantifies outcome distributions across games of varied structure and shows that LLMs draw on prior knowledge about the games and their characters – not merely the pay-off matrices – when selecting strategies.
Future work can widen the bias spectrum beyond language and personality to nationality, gender, race, age, and more; add communication among players, which might reshape payoffs and strategies [31]; and scale beyond two agents, so as to model coalition formation and commitment, as well as group reciprocity [47, 53, 48]. Such studies will clarify how trait combinations shape collective dynamics, guiding team formation, cooperative AI and socially intelligent agents [18, 36].
FAIRGAME is readily extendable to incomplete-information, simultaneous-move and sequential games, and to realistic environments where pay-offs are not pre-defined but inferred from the dynamics and outcomes of the actual use case. These scenarios enable questions such as how linguistic or personality biases influence cooperation-defection cycles in evolutionary games, or which incentives stabilize cooperation under uncertainty.
The same tool suits the rising Agentic-AI paradigm, where autonomous systems pursue complex organizational goals with minimal human oversight [4]. Modeling their interdependencies with FAIRGAME will reveal how diverse cognitive traits and contexts affect performance, supporting configurations that privilege collective benefits, over individual interests, across sectors such as healthcare, finance, manufacturing, autonomous transport, cybersecurity, smart-city infrastructures, and more.
@@
A.B. is supported by Citcom.ai, co-funded by EU/Digital Europe and, in Luxembourg, by the Feder. D.P. is supported by the European Union through the ERC INSPIRE grant (project number 101076926). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. T.A.H. is supported by EPSRC (grant EP/Y00857X/1).